OpenStudy (anonymous):

What must be the value of b so that motion of an object given by the equation D^2x+bdx+9x=0 is critically damped?

OpenStudy (anonymous):

the roots are r1 = -b+sqrt(b^2-4ac)/2a and r2 = -b-sqrt(b^2-4ac)/2a for a critically damped system, b^2 = 4ac or b^2 = 4*1*9=36 or b =6